OpenStudy (anonymous):

You find that a small business loan in the amount of 50,000 is the amount you need to purchase the restaurant location. After researching banks to find the best interest rate, you find that banks for small businesses offer the best interest rate of 9% interest that compounds monthly for 7 years. 1.What is the monthly payment for this loan? 2.Show the formula that you used and the values used for each variable to calculate the monthly payment. 3.What is the unpaid balance of the loan at the end of the 1st year? 4.Show the formula that you used and the values used for each variable to ca

OpenStudy (anonymous):

benice I can't understand your answer

OpenStudy (anonymous):

so 9% is the yearly rate... we must find the monthly rate

OpenStudy (anonymous):

\[(1+im/12)^{12}=1+i\]

OpenStudy (anonymous):

where im is the monthly interest rate and i is yearly rate

OpenStudy (anonymous):

so using a bit of math\[((1+i)^{1/12}-1)12=im\]

OpenStudy (anonymous):

0.0864878798=im

OpenStudy (anonymous):

do you understand it... my bad for the firstanswer i gave you.. its been a while..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[50k=PM \div[(1+im)^{1/12}+(1+im)^{2/12}......(1+im)^{84/12}]\]

OpenStudy (anonymous):

where pm is the monthly payment... so you have 72 payments bring the back to the present vcalue using "net present value" all of which are the same value